# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ 2.1 Causal assumptions
This assumption implies that in the pre-intervention period, the observed outcome corresponds to the potential outcome absent the policy, i.e., $\operatorname{Y}_{i,t}=\operatorname{Y}_{i,t}(0)$. Additionally, Assumption 2 precludes any anticipatory actions by economic agents, implying that they cannot alter or manipulate their outcomes prior to receiving the treatment. This assumption can be empirically tested by verifying whether pre-treatment effects are, on average, zero. Finally, the second part of Assumption 2 (post-treatment exogeneity of the covariates) is only necessary if post-treatment covariate values are used to control for post-treatment confounders (Brodersen et al.,, 2015). Although motivating covariates’ choice is often sufficient, it can be also verified by testing for the presence of treatment effects on each covariate: those that are significantly impacted by the intervention must be removed from the model.
We then make a conditional stationarity assumption (D’Haultfœuille et al.,, 2023; Hoderlein and White,, 2012).
Assumption 3
Let $\mathcal{I}_{0}$ be the information set up to time $t=0$, and denote with
$f_{\operatorname{Y}_{i,t+1}(0)|\mathcal{I}_{t},\mathbf{X}_{i,t+1}}(y_{i,t+1}(0)|\mathcal{I}_{t},\mathbf{X}_{i,t+1})$ the conditional density of the potential outcomes in the absence of the policy, where $\mathcal{I}_{t}=\{\mathcal{I}_{0},\mathbf{X}_{i,1},\dots,\mathbf{X}_{i,t},\operatorname{Y}_{i,1}(0),\dots,\operatorname{Y}_{i,t}(0)\}$. Then,
$f_{\operatorname{Y}_{i,t+1}(0)|\mathcal{I}_{t},\mathbf{X}_{i,t+1}}(y_{i,t+1}(0)|\mathcal{I}_{t},\mathbf{X}_{i,t+1})=f_{\operatorname{Y}_{i,1}(0)|\mathcal{I}_{0},\mathbf{X}_{i,1}}(y_{i,t+1}(0)|\mathcal{I}_{t},\mathbf{X}_{i,t+1}).$
Put simply, the distribution of the potential outcome absent the policy is constant for time translations, conditioned on the information set. Therefore, if we assume to know this distribution in the pre-intervention period, we would also know it in the post-intervention period. This assumption allows forecasting the path of $\operatorname{Y}_{i,t}(0)$ in the post-intervention period based on the knowledge gained from the pre-intervention period, and using it as a reliable estimator of the counterfactual potential outcome. While it is not possible to explicitly test for conditional stationarity, we can assess how well the model fits the pre-intervention data via a rigorous panel CV procedure (see Section 3). [^10] This assumption is closely related to Assumption 2. To see this, let us assume that the intervention produced an effect on the outcome before its implementation: this would inevitably alter the conditional distribution of the potential outcome, and, as a result, Assumption 3 would no longer be valid. Therefore, another way to strengthen the plausibility of this assumption is to check the forecasting performance of the assumed model by shifting the intervention date artificially backward in time. Assumption 3 implies the absence of other unforecastable shocks or policies affecting the outcome of interest in the post-intervention window. This means that the MLCM attributes any post-intervention divergence from the forecasted counterfactual only to the intervention under scrutiny.
The plausibility of this assumption hinges on institutional and contextual knowledge and requires a careful case-by-case evaluation by the researcher. Compared to popular techniques applied in the canonical setting, such as difference-in-differences and the synthetic control method, this assumption appears stronger, since those techniques can control for common shocks under the assumption that they affect treated and control units in the same way. However, in our framework, the inclusion of contemporaneous covariates plausibly unaffected by the intervention (but correlated with the outcome in the pre-intervention window) plays a similar role in that it helps controlling for changes unrelated to the treatment. [^11] To the extent that unit-specific confounders (captured by post-treatment covariates) matter more than common factors (captured by simultaneous controls in the canonical setting) as identification threats, this could make Assumption 3 more credible than the common trends assumption underlying traditional methods. This is especially true since the latter, unlike the former, must be necessarily coupled with the no-interference assumption.
More generally, the causal framework of the MLCM is characterized by a key trade-off between the plausibility of conditional stationarity (Assumption 3) and the credibility of the exogeneity of the covariate set (Assumption 2). The solution to the trade-off depends on the specific context of each application. For instance, in a study about the impact of a simultaneous but sector-specific policy change on a given outcome of interest, it might be reasonable to assume that other variables remain unaffected by the policy. Examples include a national abortion law, the EU General Data Protection Regulation, the introduction of a restrictive immigration policy, or import bans on products from a sanctioned country. For all these cases, it might be possible to come up with several predictive covariates which will likely be unaffected by the intervention at least in the short-term. In contrast, in contexts where the treatment is a large-scale, simultaneous shock with generalized impacts, such as a recession, a war, or a pandemic (as in our application, see Section 4), assuming post-treatment covariate exogeneity becomes less credible and should be avoided. [^12] Even in these cases, there may be covariates that remain unaffected, such as climate variables. The question is whether these variables are informative or helpful in controlling for potential confounders or structural breaks unrelated to the intervention. In all cases, however, it is important to acknowledge that as one moves away from the intervention date, the estimates become less credible. Indeed, not only does Assumption 3 become gradually less plausible, but post-treatment covariates may themselves become contaminated by the treatment, just as can occur with untreated units in the canonical setting (Masini and Medeiros,, 2021).
@@ 2.2 Causal estimands
In panel settings, the number of possible causal quantities increases substantially. This subsection outlines estimands that could be of general interest for observational studies based on panel data in the absence of control units. We also frame them under a finite-sample perspective, since the MLCM is thought for econometric settings in which the treatment affects, either directly or indirectly, the entire population under study. [^13] For further details on the difference between the finite-sample and superpopulation perspective in causal inference see Imbens and Rubin, (2015). In observational panel settings only Rambachan and Roth, (2023) defined causal estimands when the entire population is observed, but their approach still relies on control units. The causal estimands discussed here pertain to Case (i) defined above, namely, when the treatment affects all, or most, of the available units, simultaneously. In Appendix C, we also define causal estimands for the Case (ii) scenario with violations of the no-interference assumption. We can first define the Average Treatment Effect (ATE) across units at a given point in time.
Definition 1
For any strictly positive integer $k=1,\dots,T-t_{0}$, let $\tau_{i,t_{0}+k}=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ denote the unit-level effect of the policy at time $t_{0}+k$, where $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ is defined as the expectation of the potential outcome absent the policy given the information set, i.e., $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$ Then, the ATE at time $t_{0}+k$ is defined as,
$\tau_{t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\tau_{i,t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k|t_{0}}(0).$
@@
Note that in this benchmark scenario (Case (i)), the term $t_{0}+k$ corresponds to the Average Treatment effect on the Treated (ATT).
@@
The subsequent estimand measures the average effect within a subset of units with the same values of selected covariates. This is commonly referred to as the CATE, and it is of particular interest when there is reason to believe that the intervention has produced heterogeneous effects on different subpopulations of units as defined by their distinct characteristics. Following existing literature on CATE in high-dimensional settings with a mix of discrete and continuous covariates (Fan et al.,, 2022; Chernozhukov et al.,, 2018; Knaus et al.,, 2021), we focus on its low-dimensional summary called “group-average treatment effect”.
Definition 2
@@
Let $G_{i,t}$ denote a set of individual characteristics and indicate with $N_{g}$ the number of units in the population having $G_{i,t}=g$. For any strictly positive integer $k$, the group CATE at time $t_{0}+k$ is defined as,
$\tau_{t_{0}+k}(g)=\frac{1}{N_{g}}\sum_{i:G_{i,t_{0}+k}=g}\tau_{i,t_{0}+k}.$
@@
Note that the set of candidate conditioning variables $G_{i,t}$ used to determine CATEs does not necessarily have to match those utilized for counterfactual forecasting. This is because the variables that predict outcomes can, and often do, differ at least partially from those that predict treatment effect heterogeneity. We also remark that in a general panel setting with more than one post-treatment periods, the above definitions imply the existence of vectors representing estimated average effects, i.e., $\boldsymbol{\tau}=(\tau_{t_{0}+1},\dots,\tau_{t_{0}+k},\dots,\tau_{T})$ and $\boldsymbol{\tau(g)}=(\tau_{t_{0}+1}(g),\dots,\tau_{t_{0}+k}(g),\dots,\tau_{T}(g))$. Furthermore, sometimes researchers are also interested in temporal aggregations of such effects.
Definition 3

### v2

@@ 2.1 Assumptions
This assumption implies that in the pre-intervention period, the observed outcome corresponds to the potential outcome absent the policy, i.e., $\operatorname{Y}_{i,t}\equiv\operatorname{Y}_{i,t}(0)$ and that $\mathbf{X}_{i,t}\equiv\mathbf{X}_{i,t}(0)$ throughout all the analysis period. Additionally, Assumption 3 precludes any anticipatory actions by agents, implying that they cannot alter or manipulate their outcomes prior to receiving the treatment. This assumption can be empirically tested by verifying whether pre-treatment effects are, on average, zero. Finally, post-treatment exogeneity of the covariates is not essential for identification, as in many applications, it is more realistic to use only lagged covariates. Thus, part (iii) of Assumption 3 is relevant only in cases where post-treatment values of some covariates are needed to control for post-treatment confounders and these covariates can be considered as exogenous (Brodersen et al., 2015). [^11] Although motivating covariates’ choice is often sufficient, it can be also verified by testing for the presence of treatment effects on each covariate: those that are significantly impacted by the intervention must be removed from the model.
The next identification assumption pertains to the potential outcomes model, similar to Carvalho et al. (2018) and Masini and Medeiros (2021), but instead of using control units, we rely on a dynamic specification based on lagged outcomes.
Assumption 4 (Dynamic potential outcomes model).
Denote with $\mathbf{X}_{i,t}^{(q)}=(\mathbf{X}_{i,t},\dots,\mathbf{X}_{i,t-q})$ the collection of covariates up to time $t-q>0$, and let $\operatorname{Y}_{i,t-1}^{(p)}(0)=(\operatorname{Y}_{i,t-1}(0),\dots,\operatorname{Y}_{i,t-1-p}(0))$ be the collection of past potential outcomes up to time $t-p>0$, with $p\in\{0,\dots,t-2\}$ and $q\in\{0,\dots,t-1\}$. We assume that, for all $t=1,\dots,T$, the potential outcome absent the policy is as follows,
$\operatorname{Y}_{i,t}(0)$ $=h\left(\operatorname{Y}^{(p)}_{i,t-1}(0),\mathbf{X}^{(q)}_{i,t}\right)+\epsilon_{i,t}$ (2)
where $h(\cdot)$ is some flexible function of the past lags of the outcome, the contemporaneous covariates, and the past lags of covariates; $\epsilon_{i,t}$ is a zero-mean, uncorrelated error term following a generic distribution $f(\cdot)$
Notice that Equation (2) assumes poolability, i.e., homogeneous intercept and slope for all units, meaning that once we account for a highly predictive set of covariates and include temporal dynamics through lagged outcomes, the residual component is essentially random noise. [^12] For panel data studies with small T and large N, it is usual to pool the observations (Baltagi, 2008). We posit this assumption can be met in practice by using a subset of very predictive covariates out of a larger initial set selected ex ante on the basis of domain knowledge. This approach ensures that major sources of heterogeneity among units are effectively addressed. In addition, we can provide empirical support for this assumption by examining the fit of the pooled model during the pre-intervention period. A good fit, indicated, for instance, by low Mean Squared Error (MSE), would suggest that the pooled model adequately captures the variations in the data. In cases where the researcher also wants to account for time-invariant, group-level unobserved heterogeneity (e.g., units’ fixed effects), the covariate set could be augmented by adopting recent approaches based on sufficient representations for categorical variables (Johannemann et al., 2019).
We remark that $h(\cdot)$ can be either a linear or a non-linear function. In the linear case, no additional assumptions are needed for the identification of causal effects. In the non-linear case, no further assumptions are required if the effect is defined only a single step ahead, as in our empirical application. Conversely, identification of non-linear multi-step-ahead causal effects is notoriously difficult and has not yet been fully explored in the literature (for time series in non-causal settings, see Chen et al. (2004)). Therefore, our paper also contributes to the formal identification of causal effects in this challenging context. Specifically, when $h(\cdot)$ is non-linear and the focus is on multi-step-ahead causal effects, we propose modeling the potential outcomes in the post-intervention period as follows.
Assumption 5 (Post-intervention non-linear multi-step-ahead model).
For a positive integer $k\geq 2$, denote by $\operatorname{Y}_{t_{0}+k|t_{0}}(0)$ the expected potential outcome absent the policy, given covariates and pre-intervention lags of the outcome, i.e., $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(q)}]$. The potential outcomes absent the policy at time $t_{0}+k$ are functions of their past lags up to $t_{0}$, their conditional expectations $\operatorname{Y}_{i,t_{0}+1|t_{0}},\dots,\operatorname{Y}_{i,t_{0}+k-1|t_{0}}$, and the covariates $\mathbf{X}^{(Q)}_{i,t_{0}+k}$, with $Q=\max{\{k-1,q\}}$,
$\operatorname{Y}_{i,t_{0}+k}(0)$ $=g_{k}\left(\operatorname{Y}_{i,t_{0}+k-1|t_{0}}(0),\dots,\operatorname{Y}_{i,t_{0}+1|t_{0}}(0),\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(Q)}\right)+\epsilon_{i,t_{0}+k}$ (3)
where $\epsilon_{i,t_{0}+k}$ is a zero-mean, uncorrelated error term, and $g_{k}(\cdot)$ is a flexible function which can be different at each horizon $k$.
Notice that Equation (3) does not impose any particularly restrictive assumption on the potential outcomes. Indeed, it merely formalizes the dependence structure between potential outcomes and their conditional expectations already implied by Equation (2), e.g., for $k=2$ and $p=q=0$, we have $\operatorname{Y}_{i,t_{0}+2}(0)=h(\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)+\epsilon_{i,t_{0}+1},\mathbf{X}_{i,t_{0}+2})+\epsilon_{i,t_{0}+2}=g_{2}(\operatorname{Y}_{t_{0}+1|t_{0}}(0),\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1})+\epsilon_{i,t_{0}+2}$.
The panel CV procedure—thoroughly described in Section 3—which we propose to estimate the pre-intervention potential outcome model in Equation (2), will reveal whether the winner of the horse-race is a linear or non-linear model. If a linear model is selected, Equation (2) remains valid for all $t=1,\dots,T$. Otherwise, we can use the non-linear specification in Equation (3) to identify and estimate post-intervention counterfactual outcomes after $t_{0}+1$.
Finally, the potential outcome model defined by Equation (2) does not assume stationarity of the outcome variable. In similar contexts, studies often adopt a weaker assumption of conditional stationarity (D’Haultfœuille et al., 2023; Hoderlein and White, 2012). Formally, let $f_{\operatorname{Y}_{i,t}(0)|\operatorname{Y}_{i,t-1}^{(p)}(0),\mathbf{X}_{i,t}^{(q)}}$ be the conditional density of the potential outcomes under control. Conditional stationarity assumes that this density remains unchanged one step ahead, i.e., $f_{\operatorname{Y}_{i,t}(0)|\operatorname{Y}_{i,t-1}^{(p)}(0),\mathbf{X}_{i,t}^{(q)}}=f_{\operatorname{Y}_{i,t+1}(0)|\operatorname{Y}_{i,t}^{(p)}(0),\mathbf{X}_{i,t+1}^{(q)}}$. In other words, its functional form, which matches the form of the error term, is time-invariant, although the conditional mean can still be time-varying. [^13] Consider this simple (non-linear) example: $Y_{t}=\sin{Y_{t-1}^{2}}+\varepsilon_{t}$ where $\varepsilon_{t}\sim f(0,\sigma^{2}_{\epsilon})$. Under conditional stationarity, we have that $Y_{t}|Y_{t-1}\sim f(\sin{Y_{t-1}^{2}},\sigma^{2}_{\epsilon})$, $Y_{t+1}|Y_{t}\sim f(\sin{Y_{t}^{2}},\sigma^{2}_{\epsilon})$ and so on. Notice that while the conditional mean is time-varying, the form of the distribution $f(\cdot)$ remains the same. More generally, in our setting of interest, a short panel with small T and large N, non-stationarity is not an issue of particular concern (Baltagi, 2008). Nevertheless, as we will show below, the simulation results reported in Supplemental Appendix D demonstrate that even in the case of an explosive autoregressive coefficient (i.e., when the coefficient of the past lag of the outcome $Y_{t-1}$ is set at $\phi=1.2$), both the bias and the coverage rates of the MLCM are in line with the results obtained for stationary processes (i.e., when $\phi=0.8$). This means that the error term would be distributed in the same way during the post-intervention period. In our case, as we explain in Section 2.2, we target the causal effect at the individual level and specifically its expected value conditional on past information. Therefore, for identification, estimation, and inference, we do not require the error distribution to remain unchanged after the intervention. Instead, it is sufficient that the error term remains uncorrelated and has a mean of zero, a condition already embedded in Assumptions 4 and 5. In other words, we only require that, in the absence of the policy, the potential outcomes model would continue to be correctly specified. This implies the absence of other unforecastable shocks or policies affecting the outcome of interest in the post-intervention window. While it is not possible to explicitly test for this, we can assess how well the model fits the pre-intervention data via the rigorous panel CV procedure (see Section 3).
@@ 2.2 Causal estimands
We introduce novel estimands that could be of general interest for observational studies based on panel data in the absence of control units. The causal estimands discussed here pertain to Case (i) defined above, namely, when the treatment affects all, or most, of the available units, simultaneously. In Supplemental Appendix A, we also define causal estimands for the Case (ii) scenario with violations of the no-interference assumption.
We begin with the unconditional individual causal effect estimand.
Definition 1.
For any $k\geq 1$, the unconditional causal effect of the policy on unit $i$ at time $t_{0}+k$ is,
$\Delta_{i,t_{0}+k}=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k}(0).$ (4)
Note that this estimand arises naturally as a consequence of Assumption (2). However, identification, estimation, and inference of this causal effect would require stringent assumptions (e.g., strong stationarity of the potential outcome distribution). Therefore, our focus is on its conditional mean.
Definition 2.
For any $k\geq 1$ and some $p\in\{0,\dots,t_{0}-1\}$, $q\in\{0,\dots,t_{0}+k-1\}$ with $Q=\max{\{k-1,q\}}$, the expected individual effect of the policy, conditional on past outcomes and covariates, at time $t_{0}+k$ is,
$\tau_{i,t_{0}+k}$ $=\operatorname{E}[\left(\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k}(0)\right)|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(Q)}]$
$=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+k}^{(Q)}]$
$=\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k|t_{0}}(0).$ (5)
This estimand emphasizes that, because the potential outcomes are temporally related, the causal effect of the policy should be defined conditionally on past information. Furthermore, in many situations, the full distribution of $\Delta_{i,t_{0}+k}$ is not of interest unless we specifically target causal effects at the median or other quantiles. Typically, the mean effect is the primary concern, making $\tau_{i,t_{0}+k}$ a more suitable summary measure of policy effects. The following theorem establishes the identification conditions for the individual effect $\tau_{i,t_{0}+k}$.
Theorem 1.
For $k=1$, $\tau_{i,t_{0}+k}$ is identified under Assumptions 3 and 4. Under a linear potential outcomes model, the identifying conditions for $k\geq 2$, are the same as for $k=1$. If the potential outcomes model is non-linear, $\tau_{i,t_{0}+k}$ is identified under Assumptions 3 and 5.
Proof.
We illustrate the proof in a simple case where the potential outcome absent the policy depends only on one previous lag and contemporaneous covariates, i.e., Equations (2) and (3) become, respectively,
$\operatorname{Y}_{i,t}(0)=h(\operatorname{Y}_{i,t-1}(0),\mathbf{X}_{i,t})+\epsilon_{i,t}$
$\operatorname{Y}_{i,t_{0}+k}(0)=g_{k}(\operatorname{Y}_{i,t_{0}+k-1|t_{0}}(0),\dots,\operatorname{Y}_{i,t_{0}+1|t_{0}}(0),\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+k})+\epsilon_{i,t}$
corresponding to the case $p=q=0$. The general proof is reported in Supplemental Appendix B. The proof proceeds by induction and is organized distinguishing the two cases $k=1$ and $k=2$.
  • $k=1$. In the expression $\tau_{i,t_{0}+1}=\operatorname{Y}_{i,t_{0}+1}(1)-\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$, the first term $\operatorname{Y}_{i,t_{0}+1}(1)$ is the observed outcome under the policy and thus is immediately identified. Since we are in the case $p=q=0$, we have that $Q=\max{\{k-1,q\}}=0$, so the second term can be written as
    $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+1}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+1}]$ Def. of $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ for $p=Q=0$
    $=\operatorname{E}[h(\operatorname{Y}_{i,t_{0}}(0),\mathbf{X}_{i,t_{0}+1})|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+1}]$ Assumption 4 - Eq.(2)
    $=h(y_{i,t_{0}}(0),\mathbf{x}_{i,t_{0}+1})=y_{i,t_{0}+1|t_{0}}(0).$
    The last expression follows from the fact that $h(\cdot)$ is deterministic once we observe $\operatorname{Y}_{i,t_{0}}(0)=y_{i,t_{0}}(0)$ and $\mathbf{X}_{i,t_{0}+1}=\mathbf{x}_{i,t_{0}+1}$. As a result, $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ is identified from observed data, and so is $\tau_{i,t_{0}+1}$. Notice that in this proof, we never used the linearity of the potential outcome model. Thus, the proof at $k=1$ is the same even when $h(\cdot)$ is non-linear.
  • $k=2$. In the expression $\tau_{i,t_{0}+2}=\operatorname{Y}_{i,t_{0}+2}(1)-\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$, the first term $\operatorname{Y}_{i,t_{0}+2}(1)$ is the observed outcome under the policy. Thus, the identification of the causal effect depends solely on the second term, $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$. We now distinguish between a linear and a non-linear model specification:
    Linear case. Assume the following linear model $h(\operatorname{Y}_{i,t-1}(0),\mathbf{X_{t}})=b_{1}\operatorname{Y}_{i,t-1}(0)+\mathbf{b_{2}}\mathbf{X}_{i,t}$, where $\mathbf{b_{2}}$ is an $m-$dimensional vector of covariates’ coefficients. Since we are in the case $p=q=0$, we have that $Q=\max{\{1,0\}}=1$, so the term $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ can be written as,
    $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$ Def. $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ for $p=0,Q=1$
    $=\operatorname{E}[h(\operatorname{Y}_{i,t_{0}+1}(0),\mathbf{X}_{i,t_{0}+2})|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$ Assumption 4 - Eq.(2)
    $=\operatorname{E}[b_{1}\operatorname{Y}_{i,t_{0}+1}(0)+\mathbf{b}_{2}\mathbf{X}_{i,t_{0}+2}|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$ Linear model
    $=b_{1}\operatorname{E}[\operatorname{Y}_{i,t_{0}+1}(0)|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+1}]+\mathbf{b}_{2}\mathbf{x}_{i,t_{0}+2}$ Assumption 3
    $=b_{1}y_{i,t_{0}+1|t_{0}}(0)+\mathbf{b}_{2}\mathbf{x}_{i,t_{0}+2}.$
    Non-linear case. The identification of $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ can be based on the formulation for the post-intervention potential outcome model defined by Equation (3),
    $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}}(0),\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]\hskip 70.0pt\text{Def. $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ for $p=0,Q=1$}$
[…]
    $=\operatorname{E}[g_{2}(h(\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+1}),\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2})|\operatorname{Y}_{i,t_{0}},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
    $=g_{2}(y_{i,t_{0}+1|t_{0}}(0),y_{i,t_{0}},\mathbf{x}_{i,t_{0}+2}).$
    The last expression follows from the fact that $g_{2}(\cdot)$ is deterministic once we observe $\operatorname{Y}_{i,t_{0}}=y_{i,t_{0}}$, $\mathbf{X}_{i,t_{0}+1}=\mathbf{x}_{i,t_{0}+1}$ and $\mathbf{X}_{i,t_{0}+2}=\mathbf{x}_{i,t_{0}+2}$. In addition, even though Assumption 3 was not explicitly mentioned in this part of the proof, notice that it is implicit in Assumption 5, as by Equation (3), the potential outcomes never depend on future covariates.
The proof for a generic time $t_{0}+k$ follows analogously by induction. ∎
The following estimands are unit-averages of the individual effects defined above. For the reasons previously stated, we focus on averaging the conditional individual effects, which can be considered the panel analog of the ATE. We also frame them from a finite-sample perspective, as the MLCM is designed for statistical and econometric settings where the treatment affects, either directly or indirectly, the entire population under study. [^14] For further details on the difference between the finite-sample and superpopulation perspectives in causal inference, see Imbens and Rubin (2015). In observational panel settings, only Rambachan and Roth (2023) defined causal estimands when the entire population is observed, but their approach relies on control units.
Definition 3.
For any $k\geq 1$, the ATE at time $t_{0}+k$ is defined as the average of the individual effects across all $i$ units in the panel, with $i=1,\dots,N$,
$\tau_{t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\tau_{i,t_{0}+k}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{Y}_{i,t_{0}+k}(1)-\operatorname{Y}_{i,t_{0}+k|t_{0}}(0).$ (6)
@@
Note that in this benchmark scenario (Case (i) where only treated units are available), the term $\tau_{t_{0}+k}$ corresponds to the Average Treatment effect on the Treated (ATT).
@@
The subsequent estimand measures the average individual effect within a subset of units with the same values of selected covariates. This is commonly referred to as the CATE, and it is of particular interest when there is reason to believe that the intervention has produced heterogeneous effects on different subpopulations of units as defined by their distinct characteristics. Following existing literature on CATE in high-dimensional settings with a mix of discrete and continuous covariates (Fan et al., 2022; Knaus et al., 2021; Chernozhukov et al., 2018), we focus on its low-dimensional summary called “group-average treatment effect”.
Definition 4.
@@
Let $G_{i,t}$ denote a set of individual characteristics and indicate with $N_{g}$ the number of units in the population having $G_{i,t}=g$. For any $k\geq 1$, the CATE at time $t_{0}+k$ is defined as,
$\tau_{t_{0}+k}(g)=\frac{1}{N_{g}}\sum_{i:G_{i,t_{0}+k}=g}\tau_{i,t_{0}+k}.$ (7)
@@
Note that the set of candidate conditioning variables $G_{i,t}$ used to determine CATEs does not necessarily have to match those utilized for counterfactual forecasting. This is because the variables that predict outcomes can, and often do, differ at least partially from those that predict treatment effect heterogeneity. We also remark that in a general panel setting with more than one post-treatment periods, the above definitions imply the existence of vectors representing estimated average effects, i.e., $(\tau_{t_{0}+1},\dots,\tau_{T})$ and $(\tau_{t_{0}+1}(g),\dots,\tau_{T}(g))$. Furthermore, researchers are sometimes also interested in temporal aggregations of such effects.
Definition 5.